# StPR: Spatiotemporal Preservation and Routing for Exemplar-Free Video Class-Incremental Learning

### v1

@@ 3.4 Loss Function and Optimization
The overall loss function in our framework includes: 1) a contrastive loss between the final video feature and text description features, serving as the overall classification loss; 2) the contrast loss between the video feature of clip output under adapter fine-tuning and the text feature as the spatial level optimization; and 3) the Frame-Shared Semantics Distillation loss, which mitigates forgetting in the shared adapter modules.
Contrastive Loss Formulation. We employ a symmetric contrastive loss for both video-to-text and text-to-video alignment. Given a batch of $N$ samples, let $\mathbf{V}_{i}$ and $\mathbf{T}_{j}$ denote the video and text features, respectively. The similarity between video $i$ and text $j$ is computed as the cosine similarity $S_{i,j}=\cos(\mathbf{V}_{i},\mathbf{T}_{j})$, forming a similarity matrix $S\in\mathbb{R}^{N\times N}$. Let $\mathbf{M}\in\{0,1\}^{N\times N}$ be the label mask, where $M_{i,j}=1$ if $y_{i}=y_{j}$ and $M_{i,j}=0$ otherwise. Then, the Video-to-text contrastive loss can be caculated by:
@@
where $\varepsilon$ is a small constant added to avoid division by zero. Text-to-video contrastive loss is:
$\mathcal{L}_{\text{t2v}}=-\frac{1}{N}\sum_{j=1}^{N}\log\left(\frac{\sum_{i=1}^{N}M_{i,j}\cdot\exp(S_{i,j})}{\sum_{i=1}^{N}\exp(S_{i,j})+\varepsilon}\right).$ (14)
Symmetric total contrastive loss is :
@@
where $w$ is hyper-parameter. This design ensures that both spatial and spatiotemporal semantics are aligned with text supervision, while the FSSD term selectively preserves critical frame-shared semantics to mitigate forgetting.
@@ 4 Experiments
Table 1: Video class-incremental results on UCF101 and HMDB51, with best results in bold.
Method Exemplars Venue UCF101 HMDB51
@@
Table 2: Video class-incremental results on Kinetics400, with best results in bold.
Method Exemplars Venue Kinetics400-10s Kinetics400-20s
Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$
vCLIMB+BiC villa2022vclimb CVPR’22 $27.90$ $51.96$ $23.06$ $58.97$
vCLIMB+iCaRL villa2022vclimb CVPR’22 $32.04$ $38.74$ $26.73$ $42.25$
SMILE+BiC alssum2023just CVPR’23 $52.24$ $\mathbf{6.25}$ $48.22$ $\mathbf{0.31}$
SMILE+iCaRL alssum2023just CVPR’23 $46.58$ $7.34$ $45.77$ $4.57$
@@ 4.1 Experimental Details
Dataset. We evaluate our method on four benchmark datasets under the Video Class-Incremental Learning (VCIL) setting: UCF101 soomro2012ucf101 , HMDB51kuehne2011hmdb and Kinetics400 carreira2017quo . UCF101 and HMDB51 are widely used for human action recognition, containing 101 and 51 action classes respectively. Kinetics400 offers large-scale diversity with 400 action classes across 240,000 videos. We conduct all experiments under an exemplar-free setting. For fair comparison, we adopt the TCD protocol park2021class for UCF101 and HMDB51, where the model is pretrained on 51 and 26 base classes, and the remaining ones are distributed into 5, 10, or 25 tasks. For Kinetics400, we follow the vCLIMB benchmark villa2022vclimb , with 10- or 20-task splits.
Evaluation Metrics. We adopt two widely-used metrics to evaluate performance in VCIL: 1). Final Accuracy (Acc), which measures the overall classification accuracy on all learned classes after the final task is completed; 2). Backward Forgetting (BWF)villa2022vclimb , which quantifies the average drop in performance on previously learned tasks, reflecting how well the model retains past knowledge.
Implementation Details. All experiments are conducted on a single NVIDIA RTX 3090 GPU. We adopt the CLIP ViT-B/16 model radford2021learning as the backbone, with all its parameters frozen during training. The spatial and spatiotemporal encoders are the only trainable components in our framework. For optimization, we employ Stochastic Gradient Descent (SGD) with an initial learning rate of 0.01 and a batch size of 40. Each task is trained for 60 epochs in the first incremental session and 30 epochs in each subsequent session. The weighting hyperparameter $w$ in Eq. 16 is set to $1\times 10^{4}$. The multi-head self-attention module within the spatiotemporal encoder consists of three transformer layers, each employing two attention heads. Video clips are sampled using the TSN strategy wang2018temporal , selecting eight frames per video uniformly across the temporal dimension.
@@ 4.2 Main Results
Table. 1 and 2 report results on UCF101, HMDB51, and Kinetics400, covering different action complexities and temporal dynamics. Based on their strategies to mitigate forgetting, existing methods are categorized into two groups: 1) Exemplar-based methods.rebuffi2017icarl ; hou2019learning ; douillard2020podnet ; park2021class ; pei2022learning ; villa2022vclimb ; alssum2023just ; liang2024hypercorrelation . They store video clips, frames, or compressed features and apply rehearsal to reduce forgetting. However, these methods face scalability and privacy challenges due to their reliance on stored exemplars. 2) CIL-based methodsli2017learning ; dhar2019learning ; cheng2024stsp . This group adapts techniques from image-based class-incremental learning, such as unified distillation and subspace projection, without storing exemplars. While avoiding exemplar storage, their performance tends to be lower, especially as task difficulty increases and lacking explainable spatiotemporal disentanglement. In contrast, Our method (StPR) without storing exemplars, surpasses all baselines across datasets and settings. On UCF101 and HMDB51, it achieves 85.79% and 63.04% in the most challenging protocols, and outperforms SMILE+BiC and STSP on Kinetics400. This demonstrates the effectiveness of our core design: mitigating catastrophic forgetting by explicitly disentangling and reusing frame-shared semantics (via FSSD) and temporal dynamics (via TD-MoE) to enable robust continual learning.
@@
Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$
1 70.62 7.01 43.30 $\mathbf{6.18}$ 70.14 $\mathbf{5.33}$ 43.71 8.74
2 77.55 5.76 53.23 7.73 77.63 4.86 54.63 9.01
3 79.33 12.38 56.12 11.75 85.94 7.40 62.54 10.20
4 83.07 10.52 57.47 21.30 88.03 8.39 64.71 21.72
5 $\mathbf{85.79}$ $\mathbf{5.63}$ $\mathbf{63.04}$ 11.04 $\mathbf{88.52}$ $6.39$ $\mathbf{69.61}$ $\mathbf{7.02}$
@@ 4.3 Ablation Study
We perform ablation studies to evaluate the contribution of each component: the adapter tuning ($\mathcal{A}^{b}$), Frame-Shared Semantics Distillation (FSSD), and Temporal Decomposition-based Mixture-of-Experts (TD-MoE). Results are summarized in Table 3 and visualized across tasks in Figure 3. The baseline (pretrained CLIP) model exhibits limited performance on downstream tasks, as it lacks adaptation to new task-specific categories. Introducing FSSD alone moderately improves performance by preserving spatial semantics and reducing semantic drift, while TD-MoE independently enhances adaptation by leveraging temporal dynamics. However, using either module alone yields suboptimal performance. Combining adapter tuning with TD-MoE provides further improvements but still lacks sufficient stability in preserving spatial semantics. The full model (StPR), integrating both FSSD and TD-MoE, achieves the most stable performance across tasks. As shown in Figure 3, it maintains highest accuracy over time and effectively mitigates forgetting, demonstrating the complementary strengths of spatial semantic preservation and temporal dynamic modeling.
@@ 4.4 Further analysis
Analysis of temporal-decomposition routing strategies.
@@
Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$
Avg-MoE $81.14$ 9.95 $59.46$ 8.49 84.04 9.60 62.69 9.14
CLIP-MoE $83.59$ 7.85 $58.82$ 10.01 85.80 7.27 65.43 8.08
Adapter-MoE $83.26$ 6.07 $61.99$ $\mathbf{7.75}$ 84.31 7.82 65.57 7.68
TD-MoE(Ours) $\mathbf{85.79}$ $\mathbf{5.63}$ $\mathbf{63.04}$ 11.04 $\mathbf{88.52}$ $\mathbf{6.39}$ $\mathbf{69.61}$ $\mathbf{7.02}$
@@
Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$
w/o Distillation $83.07$ 10.52 57.47 21.30 88.03 8.39 64.71 21.72
Distillation $84.27$ 7.45 $61.74$ 13.33 88.12 7.95 67.54 13.38
FSSD(Ours) $\mathbf{85.79}$ $\mathbf{5.63}$ $\mathbf{63.04}$ $\mathbf{11.04}$ $\mathbf{88.52}$ $\mathbf{6.39}$ $\mathbf{69.61}$ $\mathbf{7.02}$
Table 4 compares our TD-MoE with several alternative Mixture-of-Experts (MoE) jacobs1991adaptive routing strategies. Simple averaging (Avg-MoE) and static weight assignments—including CLIP-MoE, which uses frozen CLIP visual features for routing, and Adapter-MoE, which uses adapter-tuned CLIP features—provide moderate improvements but fail to dynamically leverage task-specific temporal cues, often resulting in higher forgetting. In contrast, TD-MoE enables adaptive expert weighting based on temporal dynamics, consistently improving both accuracy and stability across tasks. This highlights the importance of modeling temporal variability explicitly, rather than relying on static or feature-agnostic routing.
Effectiveness of FSSD over Uniform Distillation. Table 5 compares our FSSD method with the no-distillation baseline (w/o Distillation) and standard uniform distillation (Distillation) across four VCIL settings. While uniform distillation improves accuracy and reduces backward forgetting (BWF) over the naive baseline, FSSD consistently outperforms both, achieving the highest accuracy and lowest BWF in all settings. These results highlight the benefit of selectively preserving frame-shared semantics, validating the importance-aware design of FSSD for continual video learning.
Dataset Metric $1\times 10^{3}$ $1\times 10^{4}$ $2.5\times 10^{4}$ $5\times 10^{4}$ $1\times 10^{5}$
HMDB($5\times 5$s) Acc $\uparrow$ 59.51 63.04 62.10 62.29 63.05
BWF $\downarrow$ 16.90 11.04 11.14 12.32 12.77
UCF101($5\times 10$s) Acc $\uparrow$ 85.10 85.79 84.54 84.88 84.16
BWF $\downarrow$ 7.93 5.63 6.95 5.67 6.83
Analysis of hyper-parameter. Table. 6 analyzes the sensitivity of the hyper-parameter $w$, which controls the strength of FSSD regularization. Across a wide range of values, our framework maintains stable performance, indicating robustness to hyper-parameter variations. Moderate $w$ values achieve the best trade-off between knowledge retention and adaptability, avoiding under-regularization or excessive constraint on model plasticity. For more experiments, see the appendix C.
@@ 5 Conclusion
In this work, we propose StPR, a unified and exemplar-free framework for Video Class-Incremental Learning (VCIL) to tackle the spatiotemporal challenges in continual video learning. By disentangling spatial semantics and temporal dynamics, StPR effectively mitigates catastrophic forgetting without relying on stored exemplars. Our method combines Frame-Shared Semantics Distillation (FSSD), which selectively preserves meaningful and stable semantic channels, protecting model’s plasticity. Temporal-Decomposition-based Mixture-of-Experts (TD-MoE), adaptively routes inputs based on temporal cues, reducing forgetting in deep networks. Extensive experiments on UCF101, HMDB51, and Kinetics400 validate the effectiveness and efficiency of our approach, establishing new state-of-the-art results for continual video recognition. In future work, we plan to explore more realistic application scenarios, such as open-world settings, and investigate the deployment of our method on resource-constrained edge devices.
@@ Appendix A Algorithm
A.1 Algorithm of FSSD
Input: Current task data $\mathcal{D}_{b}$; frozen model from task $b{-}1$; text features $\{\mathbf{T}_{c}\}$
@@ B.1 Frame-Shared Semantics
$\mathbb{E}[\gamma_{c,j}]\propto\mathbb{E}\left[\frac{\sum_{i}\bar{V}^{s}_{c,j}\cdot T_{c,j}}{\|\bar{\mathbf{x}}_{v}\|}\right]=\sum_{j}{\alpha_{j}\cdot\mu_{i}}\approx\frac{T_{c,j}\cdot\mu_{c,j}}{\lambda}\cdot\frac{\mu_{c,j}}{\sigma_{c,j}^{2}}.$ (25)
@@
B.2 Spatial Encoder
We adopt a frozen CLIP-ViT model enhanced with shared adapters as the spatial encoder to extract frame-level features. Let a video $V=\{V_{i}\}_{i=1}^{N_{f}}$ consist of $N_{f}$ uniformly sampled frames, where each frame $V_{i}$ is processed independently. The spatial encoder $\mathcal{F}(V_{i};\mathcal{A}^{b})$ contains $L$ transformer layers, each equipped with an adapter $\mathcal{A}_{\ell}^{b}$ inserted after the multi-head self-attention (MHSA) residual. The output of the encoder is a set of spatial features $\{\mathbf{v}_{i}^{s}\}_{i=1}^{N_{f}}$.
Transformer Block with Adapter. For the $\ell$-th transformer layer, the feature update of input token $\mathbf{v}_{i}^{(\ell)}\in\mathbb{R}^{d}$ is computed as:
$\mathbf{v}_{i}^{\prime}=\mathbf{v}_{i}^{(\ell)}+\text{MHSA}(\mathbf{v}_{i}^{(\ell)})$ (28)
$\mathbf{v}_{i}^{a}=\bm{\phi}(\mathbf{v}_{i}^{\prime\top}\mathbf{W}_{\text{down}})\mathbf{W}_{\text{up}}$ (29)
$\mathbf{v}_{i}^{\prime\prime}=\mathbf{v}_{i}^{\prime}+\mathbf{v}_{i}^{a}$ (30)
$\mathbf{v}_{i}^{(\ell+1)}=\mathbf{v}_{i}^{\prime\prime}+\text{FFN}(\mathbf{v}_{i}^{\prime\prime})$ (31)
[…]
where $\bm{\phi}(\cdot)$ is a ReLU activation, $\mathbf{W}_{\text{down}}\in\mathbb{R}^{d\times d_{h}}$ and $\mathbf{W}_{\text{up}}\in\mathbb{R}^{d_{h}\times d}$ are the adapter projection weights, $\mathbf{W}_{\text{proj}}\in\mathbb{R}^{d\times d_{vt}}$ is the projection weight that aligns the video and text feature spaces.
Multi-Head Self-Attention (MHSA). Given token sequence $\mathbf{Z}\in\mathbb{R}^{N\times d}$, multi-head self-attention is computed as:
$\text{MHSA}(\mathbf{Z})=\text{Concat}(\mathbf{h}_{1},\dots,\mathbf{h}_{H})\mathbf{W}^{O},$ (33)
where each head $\mathbf{h}_{h}$ is computed as:
$\mathbf{h}_{h}=\text{Softmax}\left(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}^{\top}}{\sqrt{d_{h}}}\right)\mathbf{V}_{h},$ (34)
with projections:
$\mathbf{Q}_{h}=\mathbf{Z}\mathbf{W}_{h}^{Q},\quad\mathbf{K}_{h}=\mathbf{Z}\mathbf{W}_{h}^{K},\quad\mathbf{V}_{h}=\mathbf{Z}\mathbf{W}_{h}^{V},$ (35)
and projection matrices $\mathbf{W}_{h}^{Q},\mathbf{W}_{h}^{K},\mathbf{W}_{h}^{V}\in\mathbb{R}^{d\times d_{h}}$, and $\mathbf{W}^{O}\in\mathbb{R}^{d\times d}$.
Feedforward Network (FFN). The FFN is a two-layer MLP with GELU activation:
$\text{FFN}(\mathbf{v})=\text{GELU}(\mathbf{v}\mathbf{W}_{1})\mathbf{W}_{2},$ (36)
where $\mathbf{W}_{1}\in\mathbb{R}^{d\times 4d}$ and $\mathbf{W}_{2}\in\mathbb{R}^{4d\times d}$.
Output. After passing through $L$ layers, the final spatial feature of the $i$-th frame is:
$\mathbf{V}_{i}^{s}=\mathcal{F}(V_{i};\mathcal{A}^{b})\in\mathbb{R}^{d_{vt}},\quad i=1,\dots,N_{f}.$ (37)
These frame-level features $\{\mathbf{V}_{i}^{s}\}_{i=1}^{N_{f}}$ are then aggregated by the spatiotemporal encoder to form the global video representation.
B.3 Spatiotemporal Encoder
Let a video clip be uniformly sampled into $N_{f}$ frames. Each frame is processed by a spatial encoder to yield frame features $\mathbf{V}^{s}_{i}$. A learnable [CLS] token $\mathbf{V}^{s}_{\text{cls}}$ is prepended to the sequence, forming the input:
$\tilde{\mathbf{V}}^{s}=[\mathbf{V}^{s}_{\text{cls}};\mathbf{V}^{s}_{1};\cdots;\mathbf{V}^{s}_{N_{f}}].$ (38)
This sequence is passed through a Transformer-based spatiotemporal encoder. Each layer comprises Multi-Head Self-Attention (MHSA), residual connections, and feedforward networks (FFNs). For a single attention head, the attention weights are computed as:
$A_{ij}=\frac{\exp\left(\frac{\langle\mathbf{q}_{i},\mathbf{k}_{j}\rangle}{\sqrt{d_{k}}}\right)}{\sum_{j^{\prime}=1}^{N_{f}+1}\exp\left(\frac{\langle\mathbf{q}_{i},\mathbf{k}_{j^{\prime}}\rangle}{\sqrt{d_{k}}}\right)},$ (39)
where $\mathbf{q}_{i}=\mathbf{W}^{Q}\tilde{\mathbf{V}}^{s}_{i}$, $\mathbf{k}_{j}=\mathbf{W}^{K}\tilde{\mathbf{V}}^{s}_{j}$ and $\mathbf{v}_{j}=\mathbf{W}^{V}\tilde{\mathbf{V}}^{s}_{j}$ are the query, key, and value projections. The output of attention for token $i$ is:
$\mathbf{y}_{i}=\sum_{j=1}^{N_{f}+1}A_{i,j}\mathbf{v}_{j}.$ (40)
We define the final spatiotemporal video representation as the output at the [CLS] token:
$\mathbf{V}^{st}=\mathbf{y}_{\text{cls}}=\sum_{j=1}^{N_{f}+1}A_{\text{cls},j}\mathbf{v}_{j}.$ (41)
This architecture enables spatiotemporal modeling by allowing: (i) long-range temporal interactions across frames via MHSA; (ii) spatial semantics retention from CLIP features; and (iii) adaptive fusion of information through the [CLS] token. Multiple attention heads further enhance expressiveness by learning diverse patterns. Consequently, $\mathbf{V}^{st}$ serves as a content-adaptive spatiotemporal descriptor capturing both motion and appearance.

### v2

@@ 3.4 Loss Function and Optimization
Our loss function includes: 1) contrastive loss between video features and text descriptions for classification; 2) contrast loss between video features under adapter fine-tuning and text features for spatial optimization; and 3) FSSD loss to mitigate forgetting in shared adapter modules.
Contrastive Loss Formulation. We use symmetric contrastive loss for video-to-text and text-to-video alignment. Given a batch of $N$ samples, let $\mathbf{V}_{i}$ and $\mathbf{T}_{j}$ denote the video and text features, respectively. The similarity between video $i$ and text $j$ is computed as the cosine similarity $S_{i,j}=\cos(\mathbf{V}_{i},\mathbf{T}_{j})$, forming a similarity matrix $S\in\mathbb{R}^{N\times N}$. Let $\mathbf{M}\in\{0,1\}^{N\times N}$ be the label mask, where $M_{i,j}=1$ if $y_{i}=y_{j}$ and $M_{i,j}=0$ otherwise. Then, the Video-to-text contrastive loss can be caculated by:
@@
where $\varepsilon$ is a small constant added to avoid division by zero. The Text-to-Video loss $\mathcal{L}_{\text{t2v}}$ is similarly defined by swapping video and text in the equation. Symmetric total contrastive loss is :
@@
where $w$ is a hyperparameter. This design aligns both spatial and spatiotemporal semantics with text supervision, while the FSSD term preserves critical frame-shared semantics to mitigate forgetting.
@@ 4 Experiments
Table 1: Average Accuracy ($\overline{\mathrm{Acc}}$) of the UCF101 and HMDB51 under the TCD benchmark.
Method Exemplar Venue UCF101 HMDB51
@@
L2P wang2022learning CVPR’22 $81.24$ $80.09$ $78.58$ $49.98$ $45.87$
S-iPrompts wang2022s NeurIPS’22 $80.60$ $80.27$ $80.43$ $53.11$ $53.89$
ST-Prompt†pei2023space CVPR’23 $84.75$ $85.54$ $85.67$ $60.14$ $60.54$
@@
StPR (Ours) $\mathbf{-}$ $\mathbf{94.67}$ $\mathbf{92.13}$ $\mathbf{88.52}$ $\mathbf{68.12}$ $\mathbf{67.01}$
Table 2: Average Accuracy ($\overline{\mathrm{Acc}}$) of the SSv2 under the TCD benchmark, with best results in bold.
Method Exemplar Venue $\textbf{10}\times\textbf{9}$s $\textbf{5}\times\textbf{18}$s
iCaRL rebuffi2017icarl CVPR’17 $20.41$ $16.62$
UCIR hou2019learning CVPR’19 $24.32$ $19.31$
PODNet douillard2020podnet ECCV’20 $27.63$ $20.14$
TCD park2021class ICCV’21 $29.32$ $24.69$
FrameMaker pei2022learning NeurIPS’22 $31.41$ $26.57$
L2P wang2022learning CVPR’22 $26.02$ $21.33$
S-iPrompts wang2022s NeurIPS’22 $33.69$ $30.84$
ST-Prompt†pei2023space CVPR’23 $39.98$ $35.44$
HCE liang2024hypercorrelation AAAI’24 $36.88$ $32.82$
StPR (Ours) $\mathbf{-}$ $\mathbf{40.79}$ $\mathbf{37.30}$
Table 3: Results of the Kinetics-400 under the vCLIMB benchmark at 10 and 20 task settings.
Method Exemplars Venue Kinetics400-10s Kinetics400-20s
@@
[…]
@@ 4.1 Experimental Details
Dataset. We evaluate our method on four benchmark datasets: UCF101 soomro2012ucf101 , HMDB51kuehne2011hmdb , Something-Something V2 (SSv2) goyal2017something and Kinetics400 carreira2017quo . All experiments are conducted in an exemplar-free setting. For fair comparison, we use the TCD benchmark park2021class on UCF101, HMDB51, and SSv2, pretraining the model on 51, 26, and 84 base classes, respectively, with the remaining classes split into tasks. For Kinetics-400, we follow the vCLIMB benchmark villa2022vclimb with 10- or 20-task splits, each containing the same number of classes.
Evaluation Metrics. We adopt three widely-used metrics to evaluate performance in VCIL: 1). Final Accuracy (Acc) villa2022vclimb , which measures the overall classification accuracy on all learned classes after the final task is completed; 2). Average Accuracy ($\overline{\mathrm{Acc}}$) park2021class , which measures the mean classification accuracy over all incremental stages after the final task is completed; 3). Backward Forgetting (BWF) villa2022vclimb , which quantifies the average drop in performance on previously learned tasks, reflecting how well the model retains past knowledge.
Implementation Details. All experiments are conducted on a single NVIDIA RTX 3090 GPU. We adopt the CLIP ViT-B/16 model radford2021learning as the backbone, with all its parameters frozen during training. The spatial and spatiotemporal encoders are the only trainable components in our framework. For optimization, we employ Stochastic Gradient Descent (SGD) with an initial learning rate of 0.01 and a batch size of 40. Each task is trained for 60 epochs in the first incremental session and 30 epochs in each subsequent session. The weighting hyperparameter $w$ in Eq. 15 is set to $1\times 10^{4}$. The multi-head self-attention module within the spatiotemporal encoder consists of three transformer layers, each employing two attention heads. Video clips are sampled using the TSN strategy wang2018temporal , selecting 8 frames per video uniformly across the temporal dimension.
@@ 4.2 Main Results
Table. 1, 2 and 3 report results on UCF101, HMDB51, SSv2 and Kinetics400, covering different action complexities and temporal dynamics. Based on their strategies to mitigate forgetting, existing methods are categorized into two groups: 1) Exemplar-based methods (iCaRL, UCIR, PODNet, TCD, FrameMaker, HCE, vCLIMB, SMILE, CSTA). They store video clips, frames, or compressed features and apply rehearsal to reduce forgetting. However, these methods face scalability and privacy challenges due to their reliance on stored exemplars. 2) CIL-based methods (LwFMC, LwM, L2P, S-iPrompts, ST-Prompt†, STSP). This group adapts techniques from image-based class-incremental learning, such as unified distillation and subspace projection, without storing exemplars. While avoiding exemplar storage, their performance tends to be lower, especially as task difficulty increases and lacking explainable spatiotemporal disentanglement. In contrast, Our method (StPR) without storing exemplars, surpasses all baselines across datasets and settings. On the TCD benchmark, our method outperforms the state-of-the-art approach (ST-Prompt†) as well as all exemplar-based methods on UCF101, HMDB51, and SSv2. On the vCLIMB benchmark, exemplar-based methods can alleviate forgetting by replaying stored samples, which makes forgetting lower. Nevertheless, our method achieves higher final accuracy, surpassing the current state-of-the-art (CSTA) and all exemplar-based counterparts.
@@
Acc $\uparrow$ $\overline{\mathrm{Acc}}\uparrow$ BWF $\downarrow$ Acc $\uparrow$ $\overline{\mathrm{Acc}}\uparrow$ BWF $\downarrow$ Acc $\uparrow$ $\overline{\mathrm{Acc}}\uparrow$ BWF $\downarrow$ Acc $\uparrow$ $\overline{\mathrm{Acc}}\uparrow$ BWF $\downarrow$
1 70.65 74.67 7.01 43.30 43.62 $\mathbf{6.18}$ 70.14 72.72 $5.33$ 43.71 47.48 8.74
2 77.55 81.84 5.76 53.23 55.67 7.73 77.63 82.06 $\mathbf{4.86}$ 54.63 60.83 9.01
3 79.33 89.36 12.38 56.12 61.14 11.75 85.94 93.47 7.40 62.54 68.88 10.20
4 83.07 91.28 10.52 57.47 63.37 21.30 88.03 94.14 8.39 64.71 73.02 21.72
5 $\mathbf{85.79}$ $\mathbf{92.13}$ $\mathbf{5.63}$ $\mathbf{63.04}$ $\mathbf{68.12}$ 11.04 $\mathbf{88.85}$ $\mathbf{94.67}$ $6.31$ $\mathbf{69.61}$ $\mathbf{75.07}$ $\mathbf{7.02}$
@@ 4.3 Ablation Study
We perform ablation studies to evaluate the contribution of each component: the adapter tuning ($\mathcal{A}^{b}$), Frame-Shared Semantics Distillation (FSSD), and Temporal Decomposition-based Mixture-of-Experts (TD-MoE). Results are summarized in Table 4. The baseline (pretrained CLIP) model exhibits limited performance on downstream tasks, as it lacks adaptation to new task-specific categories. Introducing FSSD alone moderately improves performance by preserving spatial semantics and reducing semantic drift, while TD-MoE independently enhances adaptation by leveraging temporal dynamics. However, using either module alone yields suboptimal performance. Combining adapter tuning with TD-MoE provides further improvements but still lacks sufficient stability in preserving spatial semantics. The full model (StPR), integrating both FSSD and TD-MoE, achieves the most stable performance across tasks, demonstrating the complementary strengths of spatial semantic preservation and temporal dynamic modeling.
@@ 4.4 Further analysis
Analysis of temporal-decomposition routing strategies. Table 5 compares our TD-MoE with several alternative Mixture-of-Experts (MoE) jacobs1991adaptive routing strategies. Simple averaging (Avg-MoE) and static weight assignments—including CLIP-MoE, which uses frozen CLIP visual features for routing, and Adapter-MoE, which uses adapter-tuned CLIP features—provide moderate improvements but fail to dynamically leverage task-specific temporal cues, often resulting in higher forgetting. In contrast, TD-MoE enables adaptive expert weighting based on temporal dynamics, consistently improving both accuracy and stability across tasks. This highlights the importance of modeling temporal variability explicitly, rather than relying on static or feature-agnostic routing.
@@
[…]
@@
[…]
Effectiveness of FSSD over Uniform Distillation. Table 6 compares our FSSD method with the no-distillation baseline (w/o Distillation) and standard uniform distillation (Distillation) across four VCIL settings. While uniform distillation improves accuracy and reduces backward forgetting (BWF) over the naive baseline, FSSD consistently outperforms both, achieving the highest accuracy and lowest BWF in all settings. These results highlight the benefit of selectively preserving frame-shared semantics, validating the importance-aware design of FSSD for continual video learning. For more experiments (such as hyperparameter analysis and visualization), see the appendix C
@@ 5 Conclusion
In this work, we propose StPR, a unified and exemplar-free framework for Video Class-Incremental Learning (VCIL) to tackle the spatiotemporal challenges in continual video learning. By disentangling spatial semantics and temporal dynamics, StPR effectively balances stability and plasticity without relying on stored exemplars. Our method combines Frame-Shared Semantics Distillation (FSSD), which selectively preserves meaningful and stable semantic channels, protecting model’s plasticity. Temporal-Decomposition-based Mixture-of-Experts (TD-MoE), adaptively routes inputs based on temporal cues, reducing forgetting in deep networks. Extensive experiments on UCF101, HMDB51, and SSV2 validate the effectiveness and efficiency of our approach, establishing new state-of-the-art results for continual video recognition. In future work, we plan to explore more realistic application scenarios, such as open-world settings, and investigate the deployment of our method on resource-constrained edge devices.
@@ Appendix A Algorithm
Input: Current task data $\mathcal{D}_{b}$; frozen model from task b1b{-}1; text features $\{\mathbf{T}_{c}\}$
@@ Appendix B Theoretical Supplement
$\mathbb{E}[\gamma_{c,j}]\propto\mathbb{E}\left[\frac{\sum_{i}\bar{V}^{s}_{c,j}\cdot T_{c,j}}{\|\bar{\mathbf{x}}_{v}\|}\right]=\sum_{j}{\alpha_{j}\cdot\mu_{i}}\approx\frac{T_{c,j}\cdot\mu_{c,j}}{\lambda}$ (24)